# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

### v2

@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments (RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk. Thid have been studied in detail in $d=1$ by Solomon[11] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
Does every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This is one of the key contributions of this p, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
@@
As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments. To the best of our knowledge, these theorems are novel. Our results hold for all uniformly elliptic RWRE, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
@@ Acknowledgements:
The author is an undergraduate student at the Indian Statistical Insitute, Kolkata. This work was completed during the third-year of his studies. He is deeply indebted to Dr. Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript. The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.
@@ 2. Prerequisites
Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,..,d}$ the unit vectors. Define the set
$V_{d}=\{\pm e_{i}:i=1,\dots,d\}\cup\{0\}$
@@
We call any measurable $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ as the environment function. This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$. This is essentially a mild but powerful generalisation of the environment, as we shall see.
@@
$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,\dots,d.$
@@
$\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1$
@@
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
$\mathscr{E}(\omega)(x,-v)$ $=\mathscr{E}_{i+d}(\omega)(x)\quad\text{if}\penalty 10000\ v=-e_{i}$
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{0}(\omega)\quad\text{if}\penalty 10000\ v=0.$
@@
$\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d}$
$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}$
@@
$\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}$
Therefore, the kernel of this Markov Chain are given by
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},x)$ as follows
$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
@@
$z(x+v)+z(x-v)-2z(x)<0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\}$
@@ 3. Main Results and Discussions
Let $d\geq 2$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a balanced environment function satisfying assumptions 2.2 and 2.3.
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function. Fix $\omega\in\mathcal{S}(d)$. Then under assumptions 2.2 and 2.3,
@@
converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$-$\mu$ a.e.
@@
$\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}=y}]$
@@
$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x)$
@@ 3.1. Embedded Environments
Let $d\geq 1$ and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
@@
    converges to the standard Brownian Motion on $\mathbb{R}^{d}$ with non-degenerate covariance matrix $\omega$-$\mu$ a.e.
@@ 4. An outline of the Proofs
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. This technique allows us to control the paths of the random walk which shall be key to the proof. The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
@@ 5.1. Uniqueness Principle:
    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)\quad\forall x\in D_{n}^{\circ}$.
@@
Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)\quad\forall x\in D_{n}^{\circ}$ given by
$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)]$
@@
$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in\mathbb{Z}^{d}.$
@@
$\mathbb{P}[Z_{j}^{i}=e_{i}]=\mathbb{P}_{x}[Z_{j}^{i}=-e_{i}]=\mathscr{E}_{i}(Y_{j}(0))$
@@ 5.2. Periodised Environments
We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$.
[…]
$\hat{\omega}(x)$ $=\omega([x]_{\sim}),$
$\mathscr{E}(\hat{\omega})(x)$ $=\mathscr{E}(\omega)([x]_{\sim})$
where $x\in\mathbb{Z}^{d}$. Let $\mathcal{S}_{n}$ denote the set of all such $\hat{\omega}$. Note that $\mathcal{S}_{n}\subset\mathcal{S}(d)$.
The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\hat{\omega})}$.
@@
If $g:T_{n}\to\mathbb{R}$. Define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
@@
To show that 5.5 $\implies$ 5.3, we first state a lemma by Parthsarathy[9] under the assumption that $\mu$ is stationary.
@@
[Lemma $5$ of [7]] For each $n$ there exists $\omega_{n}\in\mathcal{S}_{n}$ such that the sequence of measures $\mu_{n}$ given by
$\mu_{n}=\frac{1}{(2n)^{d}}\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}$
converges weakly to $\mu$.
@@
Proof of Theorem 5.3: Let $\{\omega_{n}\}_{n\geq 1}$ be the sequence satisfying 5.6. Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
@@
$R^{*}_{n}:l^{1}(T_{n})\to l^{\frac{d}{d-1}}(T_{n})$
Let the invariant measure for the random walk on the group $T_{n}$ given by $X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$. Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$. Therefore, one can check (or refer to Section 2 of [5]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
$\lambda_{n}(x)=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
@@
$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\penalty 10000\ \text{-a.e.}$
Since $\lambda_{n}$ extends as a measurable function over the entire $\mathbb{Z}^{d}$, we have that the family of functions
@@
$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty$
@@
$Qf(x)$ $=\mathbb{E}_{x}[\sum_{k=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(k))]$
@@
Lemma 6.2 is equivalent to the Lemma 6.1 which states that,
@@
To see the equivalence, one notes the fact that $z(x)=Qf$.
@@
$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x)$
@@
Choose $c_{3}=c_{4}\vee 1$. This completes the proof of 6.2(c).
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}$
@@
$\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}\leq\frac{d^{2}+1}{K^{2}}.$
@@
$\frac{d^{2}+1}{K^{2}}$ $<\frac{1}{2}-\frac{1}{e},\quad T_{n}\subset D_{m}$
@@
$L_{\mathscr{E}(\omega)}u(x_{0})<-f(x_{0}).$
One can choose $0<\beta<f(x_{0})$ such that
@@
It is again easy to show that.
@@
$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\omega}(k))$
@@
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x)$
@@
$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f)$
@@
$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x)$
@@
$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right)$
@@
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]$
@@
$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]$
@@
We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof and from now on we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.

### v3

@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments (RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk. This has been studied in detail in $d=1$ by Solomon[10] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
Is every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This is one of the key contributions of this paper, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
@@
As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments. To the best of our knowledge, these theorems are novel. Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
@@ Acknowledgements:
The author is an undergraduate student at the Indian Statistical Institute, Kolkata. This work was completed during the third year of his studies. He is deeply indebted to Dr. Soumendu Sundar Mukherjee of the Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript. The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.
@@ 2. Prerequisites
Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,\dots,d}$ the unit vectors. Define the set
$V_{d}=\{\pm e_{i}:i=1,\dots,d\}\cup\{0\}.$
@@
Take $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ such that $\mathscr{E}(\omega)(x)\equiv\mathscr{E}(\omega(x)):S(d)\to S(d)$ is a measurable function on $(S(d),\mathcal{B}(d))$. We call this the environment function. This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$. This is essentially a mild but powerful generalisation of the environment, as we shall see.
@@
$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega)(x)=\mathscr{E}_{i+d}(\omega)(x)\quad i=1,2,\dots,d.$
@@
$\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1.$
@@
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
$\mathscr{E}(\omega)(x,-v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{i+d}(\omega)(x)\quad\text{if}\penalty 10000\ v=-e_{i}$
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{0}(\omega)\quad\text{if}\penalty 10000\ v=0.$
@@
$\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d},$
$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}.$
@@
$\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}.$
Therefore, the kernel of this Markov Chain is given by
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Define the quantity $c(\mathscr{E},x)$ as follows
$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}.$
@@
$z(x+v)+z(x-v)-2z(x)\leq 0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\},$
@@ 3. Main Results and Discussions
Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
@@
Let $d\geq 1$. Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function. Fix $\omega\in\mathcal{S}(d)$. Then under assumptions 2.2 and 2.3,
@@
converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$ $\mu$-a.e.
@@
$G^{\mathscr{E}(\omega)}(x,y)=\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}(j)=y}],$
@@
$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x).$
@@ 3.1. Embedded Environments
Let $d\geq 1$, and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
@@
    converges to the standard Brownian Motion on $\mathbb{R}^{d}$ with non-degenerate covariance matrix $\omega$ $\mu$-a.e.
@@ 4. An outline of the Proofs
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. This technique allows us to control the paths of the random walk, which shall be key to the proof. The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
@@ 5.1. Uniqueness Principle:
    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)+u(x)\quad\forall x\in D_{n}^{\circ}$.
@@
Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)+z(x)\quad\forall x\in D_{n}^{\circ}$ given by
$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)],$
@@
$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in D_{n}.$
@@
$\mathbb{P}[Z_{j}^{i}=e_{i}]=\mathbb{P}_{x}[Z_{j}^{i}=-e_{i}]=\mathscr{E}_{i}(Y_{j})(0).$
@@ 5.2. Periodised Environments
We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathcal{S}_{n}$ denote the set of all such $\omega$.
Fix $\omega\in\mathcal{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
$\omega_{n}(x)$ $=\omega([x]_{\sim})$
$\mathscr{E}(\omega_{n})(x)$ $=\mathscr{E}(\omega)([x]_{\sim}),$
where $x\in\mathbb{Z}^{d}$.
The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\omega_{n})}$.
@@
For $g:T_{n}\to\mathbb{R}$, define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
@@
To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [5]]
@@
For each $n\geq 1$, define the sequence of measures $\mu_{n}$,
$\mu_{n}=\frac{1}{(2n)^{d}}\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}.$
Then we have:
  1. 1.
    $\mu_{n}\to\mu$ weakly $\mu$-a.e.
  2. 2.
    For every measurable $f:S(d)\to\mathbb{R}$ such that $f\in L^{p}(\mu)$ where $p\geq 1$, $\mu_{n}f\to\mu f$ $\mu$-a.e.
@@
Proof of Theorem 5.3: Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
@@
$R^{*}_{n}:l^{1}(T_{n})\to l^{\frac{d}{d-1}}(T_{n}).$
Let the invariant measure for the random walk on the group $T_{n}$ given by $\{[X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}\}_{k\geq 0}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$. Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$. Therefore, one can check (or refer to Section 2 of [4]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
$\lambda_{n}(\tau_{x}\omega_{n})=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
@@
$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\text{-a.e.}$
Since $\lambda_{n}$ extends as a measurable function over the entire $\mathcal{S}(d)$, we have that the family of functions
@@
$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty.$
@@
$Qf(x)$ $=\mathbb{E}_{x}[\sum_{k=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(k))],$
@@
Lemma 6.1 is equivalent to Lemma 6.2 which states that,
@@
To see the equivalence, one notes the fact that $z(x)=Qf(x)$.
@@
$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x).$
@@
Choose $c_{3}=c_{4}\vee 2$. This completes the proof of 6.2(c).
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
@@
$\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\sup_{x\in T_{n}}\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}=\frac{d^{2}+1}{K^{2}}.$
@@
$\frac{d^{2}+1}{K^{2}}$ $<\frac{1}{2}-\frac{1}{e},\quad T_{n}\subset D_{m}.$
@@
$L_{\mathscr{E}(\omega)}z(x_{0})<-f(x_{0})+z(x_{0}).$
One can choose $0<\beta<z(x_{0})$ such that
@@
It is again easy to show that
@@
$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
@@
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x).$
@@
$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f).$
@@
$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x).$
@@
$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right).$
@@
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})].$
@@
$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}].$
@@
We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof and from now on, we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.